OpenStudy (anonymous):

Ms. Petrova was making a green bean casserole. Her recipe called for 2 cans of the 432-gram size of green beans. She plans to triple the recipe for a holiday gathering. Approximately how many kilograms of green beans does Ms. Petrova need? A. 13 B. 1.3 C. 2.6 D. 26 Please help !

OpenStudy (apoorvk):

so each unit of the recipe requires --> 2x432 grams so 3 times the recipe will require 3 times the amount of peas per recipe=3x2x432

OpenStudy (apoorvk):

thatis in grams. so convert that into kgs.

OpenStudy (apoorvk):

what do you think would the answer be?

OpenStudy (anonymous):

and how do I do that ?

OpenStudy (apoorvk):

can you multiply 3 nos.?

OpenStudy (anonymous):

nos ?

OpenStudy (anonymous):

what do i multiply 3 by ?

OpenStudy (apoorvk):

1kg = 1000 gms so, x gms = x/1000 kgs

OpenStudy (anonymous):

so i multiply 3 x 1000 ?

OpenStudy (apoorvk):

NO!!! 3 times the recipe will require 3 times the amount of peas per recipe=3x2x432 multiply 3 by the weight of peas required per recipe.

OpenStudy (anonymous):

OHHH ; im sorry . I've never been good at math like all you smart people on here :/

OpenStudy (apoorvk):

Don't worry, none of us are smart. And neither are you bad at math. You just need to relax and go very slowly, and think. read the whole question once again, gradually, and try to understand what it asks for.

OpenStudy (anonymous):

i already multiplied 3x2x432 and got 2592 now what do i do o:

OpenStudy (apoorvk):

Wonderful! you are almost through! now that is the weight of peas in grams, right?

OpenStudy (anonymous):

correct .

OpenStudy (anonymous):

yes so how do I covert 2592 into my answer ?

OpenStudy (apoorvk):

so, if (1/1000) kg = 1g, then, 2592 g = (2592/1000) kg okay?

OpenStudy (anonymous):

2.6 ?

OpenStudy (apoorvk):

yes!

OpenStudy (anonymous):

Thank you !(:

OpenStudy (apoorvk):

Don't thank me - we are all here for the same reasons - to get help and help others. And don't fear math - it's easy. You just need to believe that you can do it and relax. :)